OpenStudy (anonymous):

(8+√5)(8-√5) A. 59-16√5 B. 64-18√5 C. 59 D.69

OpenStudy (anonymous):

Your have : \[(8+\sqrt{5})(8-\sqrt{5}) = 8\times8 - 8 \times \sqrt{5} + 8 \times \sqrt{5} - \sqrt{5} \times \sqrt{5}\] I just multiply 8 with N\[(8-\sqrt{5})\] and \[\sqrt{5}\] with \[(8-\sqrt{5})\] now it's to you to complete.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[8\times8 - 8\sqrt{5} + 8\sqrt{5} - \sqrt{5}\sqrt{5}\] = \[64 - 8\sqrt{5} + 8\sqrt{5} - \sqrt{5}\sqrt{5}\] = ?